A high-performance jet plane, practicing radar avoidance maneuvers, is in horizontal flight 40.0 m above the level ground. Suddenly, the plane encounters terrain that slopes gently upward at 3.90
Answer #1


apply from the definition of velcoity V = ds/dt

ds is the dispalcemnr and dt is time

so

dt = ds/V

Here V = 1350 * 5/18 =   375 m/s

Horizontal distance S = 40 cos 3.9 = 39.9 m

dt = 375/39.9

dt = 9.39 secs